OpenStudy (pulsified333):

Find the value(s) of cc such that the area of the region bounded by the parabolae y=x^2−c^2 and y=c^2−x^2 is 576.576. Answer (separate by commas): c=

OpenStudy (pulsified333):

@ganeshie8

ganeshie8 (ganeshie8):

start by sketching a rough diagram of the given curves and get a feel of the region whose area you seek

OpenStudy (pulsified333):

i have no clue how to do that for this

ganeshie8 (ganeshie8):

what are the given curves ?

ganeshie8 (ganeshie8):

they are simple paraboloe

ganeshie8 (ganeshie8):

y=x^2−c^2 can you sketch above parabola ?

ganeshie8 (ganeshie8):

sketch y = x^2 then pull it down by c^2 to get the graph y = x^2 - c^2

OpenStudy (pulsified333):

ok

ganeshie8 (ganeshie8):

good, similarly sketch the other curve also

ganeshie8 (ganeshie8):

y = c^2 - x^2 is same as y = -x^2 + c^2 first sketch y = -x^2, then push it up by c^2 units

OpenStudy (pulsified333):

@ganeshie8

ganeshie8 (ganeshie8):

Perfect! you seek the region between those curves

ganeshie8 (ganeshie8):

Notice that both the curves meet at their x intercepts. Can you find the x intercepts ?

OpenStudy (pulsified333):

x=-1,1

ganeshie8 (ganeshie8):

Nope, x = -c, c

ganeshie8 (ganeshie8):

looks you have taken c = 1 when you graph

OpenStudy (pulsified333):

oh ok

ganeshie8 (ganeshie8):

Can you setup a definite integral for the required area ?

OpenStudy (pulsified333):

not sure how to;(

ganeshie8 (ganeshie8):

\(\int_a^b f(x)\,dx\) gives the area between \(f(x)\) and \(x\) axis over the interval x = a and x=b.

ganeshie8 (ganeshie8):

we want to find the area between two curves over the interval x = -c and x = c

ganeshie8 (ganeshie8):

subtract the bottom curve from the top curve

ganeshie8 (ganeshie8):

Area between curves = \[\int\limits_{-c}^c \text{(top curve)}- \text{(bottom curve)}\,dx\]

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=Integrate [%28c^2-x^2%29-%28x^2-c^2%29,{x,-c,c}]

OpenStudy (pulsified333):

What do i plug in to wolfram?

ganeshie8 (ganeshie8):

plugin below in woflram Integrate[(c^2-x^2)-(x^2-c^2),{x,-c,c}]

OpenStudy (pulsified333):

(8c^3)/3

ganeshie8 (ganeshie8):

you want that to equal 576.576

ganeshie8 (ganeshie8):

(8c^3)/3 = 576.576 solve c

OpenStudy (pulsified333):

ok

OpenStudy (pulsified333):

ok got it :D